please help me to solve (c and d) knowing that d=209Question (3) 12 marks t the two primes p 41 and q 17 be given as set-up parameters for RSA a. Which of the parameters e 32,e2 49 is a valid RSA exponent? Justify your choice b. ute the private key Kpr (p, q, d). Use the extended Euclidean algorithm for the inversion and point out every calculation step. x Using the encryption key, encrypt the message M 26 X Using the private key and the Chinese remainder theory (CRT) decrypt the cipher C 513. Show all your calculation steps.
Answer #1

Given p = 41 and q = 17 from these p and q

we will get n = p * q hence n = 41 * 17 = 697

Now \Phi (n) = (p-1) * (q-1)

Hence \Phi (n) = (41-1) * (17-1) = 40 *16 = 640

a)

       We know that for public key e, gcd( e ,\Phi (n) ) = 1

       Now gcd( e1 ,\Phi (n) ) = 1 implies gcd (32,640) = 1

       this is false as gcd (32,640) != 1

       Hence for public key we will prefer e2 which is 49.

b)

      Now to find private key d we have the below relation

            e * d = 1 mod \Phi (n)

            49 * d = 1 mod 640

      Hence the value of d is 209.

c)

    For Encryption of Message M we will use the relation

   C = Me mod n   (where C is the cipher text)

Given M = 26 and we know that e = 49 and n = 697

Hence C = 2649 mod 697

   = 468

   Hence the cipher text for the plain text 26 is 468.

d)

   For Decryption of Message C we will use the relation

   M = Cd mod n   (where C is the cipher text and M is plain text)

Given C = 513 and we know that d = 209 and n = 697

   Hence M = 513209 mod 697

   = 326

   Hence the plain text for the cipher text 513 is 326.